OpenStudy (anonymous):

help please solve for x -2/7x+6+1/7x=18

OpenStudy (anonymous):

-2/7+1/7=-1/7 -1/7x+6=18 -6 -6 -1/7x=12 \[\div-1/7\] You divide -1/7 by 12 and end up with -84. So x=-84